AAVSOnet status, 2015-12-26

Sat, 12/26/2015 - 13:03

Happy holiday season, everyone!

All of the BSM systems are operational, except for BSM_Argentina (waiting for mount parts) and BSM_Hamren (waiting for camera repair).

Most of the larger telescopes (coker30, SRO, OC61) are operational.  TMO61 has had some recent dome issues that Jon Holtzman is addressing.  The azimuth drive motor wheel was replaced, and there is a pully on the slit opening mechanism that needs to be replaced to ensure that the slit will close reliably.  It may be down for another week or so, but this is full moon anyway.

Much of the recent TMO61 imaging has not been processed.  We had an issue with files not transferring in a timely manner, and had turned off the automatic nightly processing.  Matt is working on catching up, so you may see several "you've got data!" emails arriving as nights are completed.  All nights since 2015-11-18 are in this category (about 16 nights).  In addition, about a dozen nights in late October and early November are being reprocessed to pick up additional images that were not transferred in time.

We should be testing the spectrograph on TMO61 sometime during the next month, so a few nights may be spectroscopy-only.  Gary Walker is working on a possible solution to the image selector so that we can have rapid switching between imaging and spectroscopy.  A new focal reducer for the OC61 spectrograph should be on its way and should be installed during January.  We're hoping that will enable spectroscopy with the echelle to about V=7.

Software updates and reprocessing of data will be much slower for a while.  Matt is leaving as of December 31, and staff at HQ will be hard-pressed to keep up with their duties until a replacement is hired.  Please be patient during the upcoming months!